OpenStudy (anonymous):

HELP PLEASE :) If the dimensions of a rectangle are cut in half, then the new area is one fourth of the original area. True False

OpenStudy (anonymous):

If \[Area = l \times w\] We can cut both dimensions in half like this: \[\frac{ 1 }{ 2 }l \times \frac{ 1 }{ 2 }w\] Then we can multiply to find the new Area \[\frac{ 1 }{ 4 }A = \frac{ 1 }{ 2 }l \times \frac{ 1 }{ 2 }w\] So the answer is True

OpenStudy (anonymous):

thank you @Mrwingspan47

OpenStudy (anonymous):

uhmm thats not what i wrote. unless u changed your name or something. i said @Mrwingspan47

OpenStudy (anonymous):

I think it's because my name might be "inappropriate" but you're welcome :D

OpenStudy (anonymous):

i didnt write that either. ughh forget it. i wrote the name i see not the one it is showing. and i know thats why im quoting it in this text lol :)